OpenStudy (anonymous):

Radical x^2 + 21 = 2X

OpenStudy (ranga):

\[\sqrt{x^2+21} = 2x ~~ ?\]

OpenStudy (anonymous):

Yes!

OpenStudy (ranga):

Square both sides first to get rid of the radical.

OpenStudy (anonymous):

X^2 + 21 = 4x^2?

OpenStudy (ranga):

correct. subtract x^2 from both sides to isolate x.

OpenStudy (anonymous):

4x=21?

OpenStudy (ranga):

Subtract x^2 from both sides: x^2 + 21 = 4x^2 -x^2 -x^2 ------------------- ? = ? --------------------

OpenStudy (anonymous):

X+21=4?

OpenStudy (ranga):

No. You have one x-square on the left and you have four x-square on the right. If you subtract one x-square from both sides you will have: zero x-square on the left and three x-square on the right. 21 = 3x^2

OpenStudy (anonymous):

Ohh okay!

OpenStudy (ranga):

divide both sides by 3 and then take the square root on both sides to solve for x.

OpenStudy (anonymous):

2.645751311

OpenStudy (ranga):

I would leave the answer as a square root of a number rather than decimals.

OpenStudy (anonymous):

Okay thank you so much!

OpenStudy (ranga):

You are welcome. \(x = \sqrt{7}\)